Solution of (D^3 -D)y = x^3 is
1
Expert's answer
2020-03-31T15:57:28-0400

Since the indicated equation is inhomogeneous, the solution consists of two parts



"y(x)=y_{hom}(x)+y_{inhom}(x),\,\,\,\text{where}\\[0.3cm]
\left[\begin{array}{l}
\left(D^3-D\right)y=0\longrightarrow\,\,\,y=y_{hom}(x)\\[0.3cm]
\left(D^3-D\right)y=x^3\longrightarrow\,\,\,y=y_{inhom}(x),\,\,\,\text{some solution}
\end{array}\right."



1 STEP: solve a homogeneous equation



"\left(D^3-D\right)y=0\\[0.3cm]
\text{We are looking for a solution in the form}\,\,\, y(x)=e^{kx}\\[0.3cm]
Dy=\frac{d(e^{kx})}{dx}=k\cdot e^{kx}\\[0.3cm]
D^3y=\frac{d^3(e^{kx})}{dx^3}=k^3\cdot e^{kx}\\[0.3cm]
\left(D^3-D\right)y=0\longrightarrow (k^3-k)\cdot e^{kx}=0\\[0.3cm]
k\cdot(k-1)\cdot(k+1)=0\longrightarrow\left[\begin{array}{l}
k=0\\[0.3cm]
k=1\\[0.3cm]
k=-1
\end{array}\right."

Conclusion,



"\boxed{y_{hom}(x)=C_1\cdot e^{x}+C_2+C_3\cdot e^{-x}}"

2 STEP: we are looking for a particular solution to the inhomogeneous equation



"\left(D^3-D\right)y=x^3"

We are looking for a solution in the form



"y_{inhom}(x)=Ax^4+Bx^3+Cx^2+Dx+E\longrightarrow\\[0.3cm]
\left[\begin{array}{l}
Dy=\displaystyle\frac{dy_{inhom}}{dx}=4Ax^3+3Bx^2+2Cx+D\\[0.3cm]
D^3y=\displaystyle\frac{d^3y_{inhom}}{dx^3}=24Ax+6B
\end{array}\right."

Then,



"\left(D^3-D\right)y=x^3\longrightarrow\\[0.3cm]
(24Ax+6B)-(4Ax^3+3Bx^2+2Cx+D)=x^3\longrightarrow\\[0.3cm]
(-1-4A)x^3-3Bx^2+(24A-2C)x+(6B-D)=0\longrightarrow\\[0.3cm]
\left[\begin{array}{l}
x^3\,\,:\,\,-1-4A=0\longrightarrow\boxed{A=-\displaystyle\frac{1}{4}}\\[0.3cm]
x^2\,\,:\,\,-3B=0\longrightarrow\boxed{B=0}\\[0.3cm]
x^1\,\,:\,\,24A-2C=0\longrightarrow\boxed{C=-3}\\[0.3cm]
x^0\,\,:\,\,6B-D=0\longrightarrow\boxed{D=0}\\[0.3cm]
\end{array}\right."

Conclusion,



"\boxed{y_{inhom}(x)=-\frac{x^4}{4}-3x^2}"



General conclusion,


"y(x)=y_{hom}(x)+y_{inhom}(x)\longrightarrow\\[0.3cm]
\boxed{y(x)=C_1\cdot e^{x}+C_2+C_3\cdot e^{-x}-\frac{x^4}{4}-3x^2}"



ANSWER



"y(x)=C_1\cdot e^{x}+C_2+C_3\cdot e^{-x}-\frac{x^4}{4}-3x^2"